OPTICAL-OPTICAL DOUBLE RESONANCE AND LIF SPECTROSCOPY OF THE JET-COOLED BORON CARBIDE (BC) FREE RADICAL

\underline{FUMIE X. SUNAHORI}, RAMYA NAGARAJAN, AND DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

The boron carbide (^{11}BC and ^{10}BC) free radical has been produced in a pulsed discharge jet using a precursor mixture of 1% trimethylborane [B(CH_{3})_{3}] in high pressure argon. The 560 nm B^{4}Σ^{−} - X^{4}Σ^{−} electronic transition of both isotopologues has been studied in detail by LIF and emission spectroscopy. The 290 nm E^{4}Π - X^{4}Σ^{−} band system was also studied using high-resolution LIF techniques, although the rotational structure of the bands was found to be quite complicated due to significant spin-spin and spin-orbit splittings in the excited state. In order to fully assign the spectra, an optical-optical double resonance (OODR) scheme was implemented in which a yellow-green dye laser beam populated a specific rotational level of the B state and a second tunable red dye laser beam was used to promote the E^{4}Π - B^{4}Σ^{−} transition. The OODR transitions were detected by monitoring the resulting E^{4}Π - X^{4}Σ^{−} UV fluorescence. The OODR spectra consist of readily assignable spin-split P-, Q-, and R-branch lines from a single intermediate state N value, which greatly facilitated the assignment of the E - X LIF spectra. Progress in analyzing the spectra will be discussed.